# Obstructions for quantitative measure equivalence between locally compact groups

### v1

@@ 2.1 Conventions and notations
It is well known that two compact generating sets $S_{G}$ and $S^{\prime}_{G}$ of a compactly generated locally compact group $G$ provide word metrics which are bilipschitz equivalent [CDLH16, Proposition 4.B.4.(3)]. As a consequence, the volume growths associated to $S_{G}$ and $S^{\prime}_{G}$ are asymptotically equivalent (the same holds for the notions of isoperimetric profile introduced below). Thus the reason why we forget the dependency on $S_{G}$ in the notations $|.|_{G}$ and $V_{G}$ is that we only care about the asymptotics of such quantities.
@@ 2.2 Quantitative measure equivalence of locally compact groups
  • Secondly, if $\varphi$ is subadditive, being $\varphi$-integrable is the same as saying that $\int_{X}{\varphi(|c(g,x)|_{H})\mathrm{d}\nu(x)}$ is finite for all $g$ in $G$ without any uniform condition on the bound, by the same argument as in [BFS13, Appendix A.2] (see also the proof of Proposition 6.3 which yields the main ingredients). For instance, this is the case for $\varphi(t)=t^{p}$ with $0<p\leq 1$. In particular, this implies that the notion of $\varphi$-integrability does not depend on $S_{G}$ for such maps $\varphi$.
@@ 2.3 Isoperimetric profile of locally compact groups
The comparison between these notions has already been done in [Tes08, Propositions 7.1 and 7.2]. Here we provide another proof.
@@
for every $g,h\in G$. By [LNP25, Proposition 1.13], the followings are equivalent norms on $Z^{1}(G,\lambda)$:
@@ 3 A preparatory lemma
If Property (P3) is required only for the points $x$ in the intersection $X_{G}\cap X_{H}$, this lemma is relatively straightforward in the case of countable groups. Indeed, up to some translation, the fundamental domains intersect non-trivially, and it is easy to verify that $\alpha(.,x)\colon R_{X_{G}\cap X_{H}}^{G}(x)\to\alpha(R_{X_{G}\cap X_{H}}^{G}(x),x)$ and $\beta(.,x)\colon R_{X_{G}\cap X_{H}}^{H}(x)\to\beta(R_{X_{G}\cap X_{H}}^{H}(x),x)$ are bijective if $x$ lies in the intersection $X_{G}\cap X_{H}$. Since the Haar measures are the counting measures in this case, we immediately get that these bijections are measure preserving. In the non-discrete case, it is much harder and the techniques in [DLIT25, Proposition A.1] will help us, it consists in using cross-sections to slightly modify the fundamental domains so that Property (P3) in the lemma hold for every $x$ in the intersection $X_{G}\cap X_{H}$.
To get Property (P3) for almost every $x\in X_{H}$ for $\alpha$, and $x\in X_{G}$ for $\beta$, and not only for $x$ in the intersection, we require the induced actions to be ergodic, ensuring that almost every orbit in $X_{G}$ (resp. in $X_{H}$) visits the intersection $X_{G}\cap X_{H}$. To this end, we follow the proof of [KKR21, Proposition 2.17 (ii)].
Finally, we must verify that $(\varphi,\psi)$-integrability is preserved under the transformations we apply. First, for the transformations used in [KKR21, Proposition 2.17 (ii)], preservation follows from the ergodic decomposition theorem. Secondly, in [DLIT25, Proposition A.1], the new fundamental domains are obtained by translating the previous ones by elements of a compact subset of $G\times H$, ensuring a uniform bound on the difference between the length norms of the new and original cocycles.
@@
Step 1: assuming that $\mu$ is ergodic. Let us first follow the proof of [KKR21, Proposition 2.17 (ii)]. We have to be careful with the notations, since our notations $X_{H},X_{G},\mu,\nu_{G},\nu_{H},i_{G},i_{H}$ refer to the notations $X,Y,\eta,\mu,\nu,i,j$ in [KKR21]. By the Ergodic Decomposition Theorem, there is a standard probability space $(Z,\zeta)$ and a familly $((\nu_{H})_{z})_{z\in Z}$ of ergodic measures for the induced action $G\curvearrowright X_{H}$, such that
@@
for every measurable subset $A\subset X_{H}$. We then set $\mu_{z}\coloneq(i_{H})_{\star}[\lambda_{H}\otimes(\nu_{H})_{z}]$, this is a measure on $\Omega$. It is shown in the proof of [KKR21, Proposition 2.17 (ii)] that the measures $\mu_{z}$ are $\sigma$-finite and ergodic (with respect to the $(G\times H)$-action on $\Omega$) for almost all $z\in Z$, and give rise to measures $(\nu_{G})_{z}$ on $X_{G}$ such that
@@
Step 2: Property (P3) for $x$ in the intersection $X_{G}\cap X_{H}$. Let us fix $z$ in a conull subset of $Z$ for which the desired properties hold. By [DLIT25, Proposition A.1], we can choose new fundamental domains $X^{\prime}_{G,z}$ and $X^{\prime}_{H,z}$ satisfying Property (P1), and so that the new cocycles $\alpha^{\prime}_{z}$ and $\beta^{\prime}_{z}$ satisfy Property (P3) for every $x\in X^{\prime}_{G,z}\cap X^{\prime}_{H,z}$. Furthermore, these cocycles are $\varphi$- and $\psi$-integrable, since their norms are bounded above by the norms of the previous cocycles with an additive term (see the proof of [DLIT25, Proposition A.1]).
@@
are measure preserving, and the goal is now to extend it to $y\in X_{H}$ (for $\alpha_{z}^{\prime}$) and to $y\in X_{G}$ (for $\beta_{z}^{\prime}$), using ergodicity. Coming back to Step 1 of this proof, or equivalently to the proof of [KKR21, Proposition 2.17 (ii)], we know that the ergodic measure $\mu_{z}$ gives rise to ergodic finite measures $(\nu^{\prime}_{H})_{z}$ and $(\nu^{\prime}_{G})_{z}$ for the induced actions $G\curvearrowright X^{\prime}_{H,z}$ and $H\curvearrowright X^{\prime}_{G,z}$. To conclude, for almost every $x\in X^{\prime}_{H,z}$, the set $R_{X^{\prime}_{G,z}\cap X^{\prime}_{H,z}}^{G}(x)$ is not empty and we pick a point $g_{x}$. Then for every $g\in R_{X^{\prime}_{G,z}\cap X^{\prime}_{H,z}}^{G}(x)$, the cocycle identity gives
@@
where $g_{x}\cdot x$ lies in the intersection $X^{\prime}_{G,z}\cap X^{\prime}_{H,z}$, so it follows from unimodularity and from the fact that the map (8) (with $y=g_{x}\cdot x$) is measure preserving, that
@@ 4 Behaviour of volume growth
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups and $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a non-decreasing and subadditive map. If there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
We begin with a simple result, found for instance in [Aus16, Proof of Lemma B.11].
@@
Let $G\curvearrowright(X,\nu)$ be a pmp action of a locally compact group on a probability space. Let $Y$ be a measurable subset of $X$, of positive measure. Then we have
@@
Without loss of generality, we assume that the properties given by Lemma 3.1 hold. Let us also assume that $\nu_{H}(X_{H})=1$. Let $Y\coloneq X_{G}\cap X_{H}$. Let $K,K^{\prime}>0$ be positive constants such that
@@
By [DKLMT22, Lemma 2.24] which also holds in our framework, there exists a constant $C>0$ such that for every $g\in G$,
@@ 5 Behaviour of isoperimetric profiles
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups and $p\geq 1$. If there exists an $(\mathrm{L}^{p},\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
Let $G$ and $H$ be non-discrete unimodular locally compact compactly generated groups and $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a non-decreasing map such that $t\mapsto t/\varphi(t)$ is non-decreasing. If there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@ 5.1 Outlines of the proofs
As for Theorem A, we assume without loss of generality that the properties given by Lemma 3.1 hold.
@@ 5.2 Proof of the theorems
We follow the steps described in Section 5.1 to prove Theorems 5.1 and 5.2. In the following propositions and lemmas, we assume that the properties given by Lemma 3.1 hold, and we fix a map $f\colon H\to\mathbb{R}$ giving rise to a random map $f_{x}\colon G\to\mathbb{R}$ as described above. Moreover $G$ and $H$ are always assumed to be non-discrete unimodular locally compact compactly generated groups.
@@
We assume without loss of generality that the measure equivalence coupling satisfies the properties provided by Lemma 3.1. Let $v$ be a positive real number, let $f\colon H\to\mathbb{R}$ whose support has measure less than $v$ and the associated maps $f_{x}$ as explained in Section 5.1. Let $K,K^{\prime},C_{p}$ be the constants introduced in Propositions 5.3, 5.4 and 5.8. Note that $C_{p}$ is finite since the cocycle is $\mathrm{L}^{p}$ and by the comments after Definition 2.4 (namely $S_{G}$ can be replaced by $S_{G}^{-1}$).
@@
We assume without loss of generality that the measure equivalence coupling satisfies the properties provided by Lemma 3.1. Let $v$ be a positive real number, let $f\colon H\to\mathbb{R}$ whose support has measure less than $v$ and the associated maps $f_{x}$ as explained in Section 5.1. We assume that $\|\nabla^{\sup}_{H}f\|_{1}=1$ (it is easy to see that the definition of the isoperimetric profile can be restricted to such functions). Let $K,K^{\prime},C_{\varphi}$ be the constants introduced in Propositions 5.3, 5.4 and 5.9. Since we have $\|\nabla^{\sup}_{H}f\|_{1}=1$, the constant $C_{\varphi}$ does not depend on $f$. Moreover, this constant is finite by assumption and the comments after Definition 2.4 (namely $S_{G}$ can be replaced by $S_{G}^{-1}$).

### v2

@@ 2.1 Conventions and notations
It is well known that two compact generating sets $S_{G}$ and $S^{\prime}_{G}$ of a compactly generated locally compact group $G$ provide word metrics which are bilipschitz equivalent [5, Proposition 4.B.4.(3)]. As a consequence, the volume growths associated to $S_{G}$ and $S^{\prime}_{G}$ are asymptotically equivalent (the same holds for the notions of isoperimetric profile introduced below). Thus the reason why we forget the dependency on $S_{G}$ in the notations $|.|_{G}$ and $V_{G}$ is that we only care about the asymptotics of such quantities.
@@ 2.2 Quantitative measure equivalence of locally compact groups
  • Secondly, if $\varphi$ is subadditive, being $\varphi$-integrable is the same as saying that $\int_{X}{\varphi(|c(g,x)|_{H})\mathrm{d}\nu(x)}$ is finite for all $g$ in $G$ without any uniform condition on the bound, by the same argument as in [2, Appendix A.2] (see also the proof of Proposition 6.3 which yields the main ingredients). For instance, this is the case for $\varphi(t)=t^{p}$ with $0<p\leq 1$. In particular, this implies that the notion of $\varphi$-integrability does not depend on $S_{G}$ for such maps $\varphi$.
@@
Finally, since we will make use of [10, Proposition A.1] which requires non-discrete groups, we will have to reduce the problem to this setting. The following will be particularly useful.
Lemma 2.6.
Let $G$ and $H$ be unimodular locally compact compactly generated groups, and $\varphi,\psi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be increasing maps satisfying $\varphi(Ct)=O(\varphi(t))$ and $\psi(Ct)=O(\psi(t))$ for every constant $C>0$. Given non-discrete compact groups $K_{G}$ and $K_{H}$, let $\tilde{G}\coloneq G\times K_{G}$ and $\tilde{H}\coloneq K\times K_{H}$. Then the following properties hold:
  • any $(\varphi,\psi)$-integrable measure equivalence coupling from $G$ and $H$ gives rise to $(\varphi,\psi)$-integrable measure equivalence coupling from $\tilde{G}$ and $\tilde{H}$;
  • for every $p\geq 1$, $j_{p,G}(n)\simeq j_{p,\tilde{G}}(n)$ and $j_{p,H}(n)\simeq j_{p,\tilde{H}}(n)$;
  • $V_{G}(n)\simeq V_{\tilde{G}}(n)$ and $V_{H}(n)\simeq V_{\tilde{H}}(n)$.
In other words, in Theorems A, B, C, D and E, we can assume without loss of generality that the groups $G$ and $H$ are not discrete.
Proof.
Invariance of isoperimetric profiles and volume growth follows from the fact that $G$ and $\tilde{G}$ are quasi-isometric, as well as $H$ and $\tilde{H}$. Note that, by compactness, the Haar measures $\lambda_{K_{G}}$ and $\lambda_{K_{H}}$ of $K_{G}$ and $K_{H}$ are finite, this is a crucial point for the sequel. Let $(\Omega,X_{G},X_{H},\mu)$ be a $(\varphi,\psi)$-integrable measure equivalence coupling from $G$ to $H$, with the maps $i_{G}$ and $i_{H}$ as in Definition 2.1 and the cocycles $\alpha\colon G\times X_{H}\to H$ and $\beta\colon H\times X_{G}\to G$ as in Definition 2.2. A coupling between $\tilde{G}$ and $\tilde{H}$ is built as follows:
  • $\tilde{\Omega}=\Omega\times K_{G}\times K_{H}$, $\tilde{\mu}=\mu\otimes\lambda_{K_{G}}\otimes\lambda_{K_{H}}$;
  • $\tilde{X}_{\tilde{G}}=X_{G}\times K_{H}$, $\tilde{\nu}_{\tilde{G}}=\nu_{G}\otimes\lambda_{K_{H}}$;
  • $\tilde{X}_{\tilde{H}}=X_{H}\times K_{G}$, $\tilde{\nu}_{\tilde{H}}=\nu_{H}\otimes\lambda_{K_{G}}$;
  • $\tilde{\imath}_{\tilde{G}}\colon((g,k_{G}),(x_{G},k_{H}))\in\tilde{G}\times\tilde{X}_{\tilde{G}}\mapsto(i_{G}(g,x_{G}),k_{G},k_{H})\in\Omega$;
  • $\tilde{\imath}_{\tilde{H}}\colon((h,k_{H}),(x_{H},k_{G}))\in\tilde{H}\times\tilde{X}_{\tilde{H}}\mapsto(i_{H}(h,x_{H}),k_{G},k_{H})\in\Omega$,
[…]
$(g,k_{G})\ast\tilde{\imath}_{G}((g^{\prime},k^{\prime}_{G}),(x_{G},k_{H}))$ $=(i_{G}(gg^{\prime},x_{G}),k_{G}k^{\prime}_{G},k_{H})$
$=(g\ast i_{G}(g^{\prime},x_{G}),k_{G}k^{\prime}_{G},k_{H}),$
$(h,k_{H})\ast\tilde{\imath}_{H}((h^{\prime},k^{\prime}_{H}),(x_{H},k_{G}))$ $=(i_{H}(hh^{\prime},x_{H}),k_{G},k_{H}k^{\prime}_{H})$
$=(h\ast i_{H}(h^{\prime},x_{H}),k_{G},k_{H}k^{\prime}_{H}).$
It is straightforward to prove that it defines a measure equivalence coupling. Moreover, the cocycles $\tilde{\alpha}\colon\tilde{G}\times\tilde{X}_{\tilde{H}}\to\tilde{H}$ and $\beta\colon\tilde{H}\times\tilde{X}_{\tilde{G}}\to\tilde{G}$ are defined by:
$\tilde{\alpha}((g,k_{G}),(x_{H},k^{\prime}_{G}))=(\alpha(g,x_{H}),1_{K_{G}}),$
$\tilde{\beta}((h,k_{H}),(x_{G},k^{\prime}_{H}))=(\beta(h,x_{G}),1_{K_{H}}),$
this immediately implies that $\alpha$ is $\varphi$-integrable and $\beta$ is $\psi$-integrable. ∎
@@ 2.3 Isoperimetric profile of locally compact groups
The comparison between these notions has already been done in [30, Propositions 7.1 and 7.2]. Here we provide another proof.
@@
for every $g,h\in G$. By [24, Proposition 1.13], the following are equivalent norms on $Z^{1}(G,\lambda)$:
@@ 3 A preparatory lemma
If Property (P3) is required only for the points $x$ in the intersection $X_{G}\cap X_{H}$, this lemma is relatively straightforward in the case of countable groups. Indeed, up to some translation, the fundamental domains intersect non-trivially, and it is easy to verify that $\alpha(.,x)\colon R_{X_{G}\cap X_{H}}^{G}(x)\to\alpha(R_{X_{G}\cap X_{H}}^{G}(x),x)$ and $\beta(.,x)\colon R_{X_{G}\cap X_{H}}^{H}(x)\to\beta(R_{X_{G}\cap X_{H}}^{H}(x),x)$ are bijective if $x$ lies in the intersection $X_{G}\cap X_{H}$. Since the Haar measures are the counting measures in this case, we immediately get that these bijections are measure preserving. In the non-discrete case, it is much harder and the techniques in [10, Proposition A.1] will help us, it consists in using cross-sections to slightly modify the fundamental domains so that Property (P3) in the lemma hold for every $x$ in the intersection $X_{G}\cap X_{H}$.
To get Property (P3) for almost every $x\in X_{H}$ for $\alpha$, and $x\in X_{G}$ for $\beta$, and not only for $x$ in the intersection, we require the induced actions to be ergodic, ensuring that almost every orbit in $X_{G}$ (resp. in $X_{H}$) visits the intersection $X_{G}\cap X_{H}$. To this end, we follow the proof of [22, Proposition 2.17 (ii)].
Finally, we must verify that $(\varphi,\psi)$-integrability is preserved under the transformations we apply. First, for the transformations used in [22, Proposition 2.17 (ii)], preservation follows from the ergodic decomposition theorem. Secondly, in [10, Proposition A.1], the new fundamental domains are obtained by translating the previous ones by elements of a compact subset of $G\times H$, ensuring a uniform bound on the difference between the length norms of the new and original cocycles.
@@
Step 1: assuming that $\mu$ is ergodic. Let us first follow the proof of [22, Proposition 2.17 (ii)]. We have to be careful with the notations, since our notations $X_{H},X_{G},\mu,\nu_{G},\nu_{H},i_{G},i_{H}$ refer to the notations $X,Y,\eta,\mu,\nu,i,j$ in [22]. By the Ergodic Decomposition Theorem, there is a standard probability space $(Z,\zeta)$ and a familly $((\nu_{H})_{z})_{z\in Z}$ of ergodic measures for the induced action $G\curvearrowright X_{H}$, such that
@@
for every measurable subset $A\subset X_{H}$. We then set $\mu_{z}\coloneq(i_{H})_{\star}[\lambda_{H}\otimes(\nu_{H})_{z}]$, this is a measure on $\Omega$. It is shown in the proof of [22, Proposition 2.17 (ii)] that the measures $\mu_{z}$ are $\sigma$-finite and ergodic (with respect to the $(G\times H)$-action on $\Omega$) for almost all $z\in Z$, and give rise to measures $(\nu_{G})_{z}$ on $X_{G}$ such that
@@
Step 2: Property (P3) for $x$ in the intersection $X_{G}\cap X_{H}$. Let us fix $z$ in a conull subset of $Z$ for which the desired properties hold. By [10, Proposition A.1], we can choose new fundamental domains $X^{\prime}_{G,z}$ and $X^{\prime}_{H,z}$ satisfying Property (P1), and so that the new cocycles $\alpha^{\prime}_{z}$ and $\beta^{\prime}_{z}$ satisfy Property (P3) for every $x\in X^{\prime}_{G,z}\cap X^{\prime}_{H,z}$. Furthermore, these cocycles are $\varphi$- and $\psi$-integrable, since their norms are bounded above by the norms of the previous cocycles with an additive term (see the proof of [10, Proposition A.1]).
@@
are measure preserving, and the goal is now to extend it to $y\in X_{H}$ (for $\alpha_{z}^{\prime}$) and to $y\in X_{G}$ (for $\beta_{z}^{\prime}$), using ergodicity. Coming back to Step 1 of this proof, or equivalently to the proof of [22, Proposition 2.17 (ii)], we know that the ergodic measure $\mu_{z}$ gives rise to ergodic finite measures $(\nu^{\prime}_{H})_{z}$ and $(\nu^{\prime}_{G})_{z}$ for the induced actions $G\curvearrowright X^{\prime}_{H,z}$ and $H\curvearrowright X^{\prime}_{G,z}$. To conclude, for almost every $x\in X^{\prime}_{H,z}$, the set $R_{X^{\prime}_{G,z}\cap X^{\prime}_{H,z}}^{G}(x)$ is not empty and we pick a point $g_{x}$. Then for every $g\in R_{X^{\prime}_{G,z}\cap X^{\prime}_{H,z}}^{G}(x)$, the cocycle identity gives
@@
where $g_{x}\cdot x$ lies in the intersection $X^{\prime}_{G,z}\cap X^{\prime}_{H,z}$. It finally follows from unimodularity, the equality $\left(R_{X^{\prime}_{G,z}\cap X^{\prime}_{H,z}}^{G}(x)\right)g_{x}^{-1}=R_{X^{\prime}_{G,z}\cap X^{\prime}_{H,z}}^{G}(g_{x}\cdot x)$ and the fact that the map (8) (with $y=g_{x}\cdot x$) is measure preserving, that
@@ 4 Behaviour of volume growth
Let $G$ and $H$ be unimodular locally compact compactly generated groups and $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a non-decreasing and subadditive map. If there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
We begin with a simple result, found for instance in [1, Proof of Lemma B.11].
@@
Let $G\curvearrowright(X,\nu)$ be a measure preserving action of a locally compact group on a probability space. Let $Y$ be a measurable subset of $X$, of positive measure. Then we have
@@
Without loss of generality, we assume that the groups are non-discrete (using Lemma 2.6) and that the properties given by Lemma 3.1 hold. Let us also assume that $\nu_{H}(X_{H})=1$. Let $Y\coloneq X_{G}\cap X_{H}$. Let $K,K^{\prime}>0$ be positive constants such that
@@
By [9, Lemma 2.24] which also holds in our framework, there exists a constant $C>0$ such that for every $g\in G$,
@@ 5 Behaviour of isoperimetric profiles
Let $G$ and $H$ be unimodular locally compact compactly generated groups and $p\geq 1$. If there exists an $(\mathrm{L}^{p},\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@
Let $G$ and $H$ be unimodular locally compact compactly generated groups and $\varphi\colon\mathbb{R}_{+}\to\mathbb{R}_{+}$ be a non-decreasing map such that $t\mapsto t/\varphi(t)$ is non-decreasing. If there exists a $(\varphi,\mathrm{L}^{0})$-integrable measure equivalence coupling from $G$ to $H$, then
@@ 5.1 Outlines of the proofs
As for Theorem A, we assume without loss of generality that the groups are non-discrete (using Lemma 2.6) and the properties given by Lemma 3.1 hold.
@@ 5.2 Proof of the theorems
We follow the steps described in Section 5.1 to prove Theorems 5.1 and 5.2. In the following propositions and lemmas, we assume that the groups are non-discrete (using Lemma 2.6) and the properties given by Lemma 3.1 hold, and we fix a map $f\colon H\to\mathbb{R}$ giving rise to a random map $f_{x}\colon G\to\mathbb{R}$ as described above.
@@
We assume without loss of generality that the groups are non-discrete (using Lemma 2.6) and that the measure equivalence coupling satisfies the properties provided by Lemma 3.1. Let $v$ be a positive real number, let $f\colon H\to\mathbb{R}$ whose support has measure less than $v$ and the associated maps $f_{x}$ as explained in Section 5.1. Let $K,K^{\prime},C_{p}$ be the constants introduced in Propositions 5.3, 5.4 and 5.8. Note that $C_{p}$ is finite since the cocycle is $\mathrm{L}^{p}$ and by the comments after Definition 2.4 (namely $S_{G}$ can be replaced by $S_{G}^{-1}$).
@@
We assume without loss of generality that the groups are non-discrete (using Lemma 2.6) and the measure equivalence coupling satisfies the properties provided by Lemma 3.1. Let $v$ be a positive real number, let $f\colon H\to\mathbb{R}$ whose support has measure less than $v$ and the associated maps $f_{x}$ as explained in Section 5.1. We assume that $\|\nabla^{\sup}_{H}f\|_{1}=1$ (it is easy to see that the definition of the isoperimetric profile can be restricted to such functions). Let $K,K^{\prime},C_{\varphi}$ be the constants introduced in Propositions 5.3, 5.4 and 5.9. Since we have $\|\nabla^{\sup}_{H}f\|_{1}=1$, the constant $C_{\varphi}$ does not depend on $f$. Moreover, this constant is finite by assumption and the comments after Definition 2.4 (namely $S_{G}$ can be replaced by $S_{G}^{-1}$).